OpenStudy (anonymous):

Find the distance driven, if you drive at 60 mph for three and a half hours. Round to the nearest whole mile. 17 mi 180 mi 210 mi 215 mi

OpenStudy (anonymous):

please dont put physics problems in chemistry. v=s/t -> s= v*t!!!!!

OpenStudy (anonymous):

answer 210 mi